# EquiAV: Leveraging Equivariance for Audio-Visual Contrastive Learning

### v1

@@ Abstract
Recent advancements in self-supervised audio-visual representation learning have demonstrated its potential to capture rich and comprehensive representations. However, despite the advantages of data augmentation verified in many learning methods, audio-visual learning has struggled to fully harness these benefits, as augmentations can easily disrupt the correspondence between input pairs. To address this limitation, we introduce EquiAV, a novel framework that leverages equivariance for audio-visual contrastive learning. Our approach begins with extending equivariance to audio-visual learning, facilitated by a shared attention-based transformation predictor. It enables the aggregation of features from diverse augmentations into a representative embedding, providing robust supervision. Notably, this is achieved with minimal computational overhead. Extensive ablation studies and qualitative results verify the effectiveness of our method. EquiAV outperforms previous works across various audio-visual benchmarks.
Machine Learning
@@ 1 Introduction
Audio and visual modalities play a pivotal role in how humans perceive their surroundings. Despite the differences in their characteristics, there exists an inherent correspondence between the two modalities. Learning such audio-visual correspondence from large-scale unlabeled video data in a self-supervised manner has recently become a major interest in the deep-learning research community. Among various approaches for audio-visual self-supervised learning, Audio-Visual Contrastive Learning has been popular due to its simplicity and effectiveness (Ma et al., 2021a, b; Morgado et al., 2021a, b; Patrick et al., 2021; Recasens et al., 2021; Wang et al., 2021; Sarkar & Etemad, 2023).
@@
A key challenge in audio-visual contrastive learning is to enhance representational capability and diversity while maintaining the correspondence between two different modalities. The most straightforward and effective way to enrich the representation is to utilize data augmentation. However, data augmentation has not been extensively applied in the field of audio-visual contrastive learning. This is because correspondence between audio and visual modalities can be easily distorted due to data augmentations. For this reason, previous works utilize a very limited range of data augmentation and adopt alternatives to learn richer representations, such as incorporating masked data modeling (Gong et al., 2023; Huang et al., 2023) or utilizing temporal supervision (Sarkar & Etemad, 2023; Jenni et al., 2023).
@@ 2 Related Works
Audio-visual contrastive learning has been one of the most popular approaches for learning the natural correspondence between audio and visual modalities, due to its simple intuition and powerful performance on the downstream tasks (Ma et al., 2021a, b; Morgado et al., 2021a, b; Patrick et al., 2021; Recasens et al., 2021; Wang et al., 2021; Owens & Efros, 2018; Sarkar & Etemad, 2023). A common approach to audio-visual contrastive learning is to learn the context of the synchronous relationship between audio and visual inputs (Korbar et al., 2016; Alwassel et al., 2020; Morgado et al., 2021b; Sarkar & Etemad, 2023). On the other hand, several works (Georgescu et al., 2023; Haliassos et al., 2022) adopt masked modeling techniques that are designed to reconstruct the masked raw inputs or predict the masked context features. Recently, CAV-MAE (Gong et al., 2023) and MAViL (Huang et al., 2023) have combined contrastive learning and masked data modeling techniques to learn complementary representations.
@@ 3.2 EquiAV
In our framework, typical visual-based augmentations are employed for visual inputs. However, unlike the abundant augmentations available for an image or video, the options for the audio modality are relatively limited. To address this issue, we convert the audio signal into a spectrogram, followed by the application of visual-based augmentations in the same manner. The augmentation information is encoded into real vectors, denoted as $\mathbf{t}_{a}$ and $\mathbf{t}_{v}$. These augmentation vectors parameterize how much each augmentation is applied to the input data. A detailed explanation of data augmentation parameterization is provided in Appendix B.
@@
$\bar{h}=\mathbb{E}_{h^{\prime}\sim D_{h}}\left[h^{\prime}\right].$ (11)
The vector $\bar{h}$ integrates information from multiple augmented data, securing more general features while also providing robustness against undesirable effects of augmentation. However, it is not feasible to generate all transformations and obtain the representative vector $\bar{h}$, we aim to approximate it with the centroid of sampled equivariant representations.
$\displaystyle\bar{h}\approx\bar{h}_{m}$ $=\frac{1}{S}\sum_{i=1}^{S}\hat{h}_{m}^{i}=\frac{1}{S}\sum_{i=1}^{S}u_{m}(h_{m}^{i},\mathbf{t}_{m}^{i}).$ (12)
@@
    $\mathcal{L}^{\text{inter}}=\frac{1}{N}\sum_{k=1}^{N}\ell^{\text{inter}}(k)$
@@ 4 Experiments
Table 1: Audio-visual event classification performance on AudioSet and VGGSound. A: Audio-only, V: Visual-only, A-V: Audio-visual. IN SL: ImageNet supervised learning, SSL: Self-supervised learning, $*{}^{*}$ Non-standard train/test split.
Pretrain AudioSet-20K (mAP) AudioSet-2M (mAP) VGGSound (acc.)
Method A V A-V A V A-V A V A-V
GBlend (Wang et al., 2020) - 29.1 22.1 37.8 32.4 18.8 41.8 - - -
Perceiver (Jaegle et al., 2021) - - - - 38.4 25.8 44.2 - - -
Attn AV (Fayek & Kumar, 2020) IN SL - - - 38.4 25.7 46.2 - - -
$*{}^{*}$ (Nagrani et al., 2021) IN21K SL 31.3 27.7 43.9 41.5 31.3 49.6 52.3 51.2 64.1
CAV-MAE (Gong et al., 2023) SSL 37.7 19.8 42.0 46.6 26.2 51.2 59.5 47.0 65.5
$*{}^{*}$ (Georgescu et al., 2023) SSL - - - 46.6 31.1 51.8 57.2 50.3 65.0
MAViL (Huang et al., 2023) SSL 41.8 24.8 44.9 48.7 30.3 53.3 60.8 50.9 67.1
EquiAV (ours) SSL 42.4 25.7 46.6 49.1 30.1 54.6 61.0 50.7 67.1
The model is pre-trained on the AudioSet-2M in a self-supervised manner, without the use of labels. The audio encoder $f_{a}$ and visual encoder $f_{v}$ are initialized with the self-supervised pre-trained ViT-B/16 model of MAE (He et al., 2022). We employ 3-layer MLP with layer normalization for intra-modal and inter-modal projection heads. Note that the multi-head attention layer and the feed-forward layer of the transformation predictor are similar to those used in Transformers (Vaswani et al., 2017; Chen et al., 2021), which include layer normalization layer and residual connection. We set $\lambda^{\text{inter}}$, $\lambda^{\text{inter}}_{a}$ and $\lambda^{\text{inter}}_{v}$ all the same as 1. More detailed experimental settings are explained in the Appendix B and Appendix C.
Table 2: Audio classification accuracy on ESC-50 and SPC-v1. ${}^{\dagger}$esults reproduced on our environment. [^2] Weights from https://github.com/YuanGongND/cav-mae
Method ESC-50 SPC-v1
${}^{\dagger}$ (Gong et al., 2023) 84.0 96.4
${}^{\dagger}$ (Gong et al., 2023) 83.2 97.1
MAViL (Huang et al., 2023) 94.4 97.4
EquiAV (ours) 96.0 97.8
@@ 4.1 Main Results
Fine-Tuning on the Single-modal Benchmarks.
For the single-modal downstream tasks, we fine-tune the audio branch of our model with the Environmental Sound Classification (ESC-50) (Piczak, 2015) and the Speech Commands (SPC-v1) (Warden, 2018) datasets. According to the results reported in Table 2, our method shows superior performance compared to other self-supervised learning-based methods.
@@
Table 3 reports the performance of zero-shot retrieval on audio-visual datasets, based on the similarity between the audio-visual embeddings. The sample lists used in the experiments for AudioSet and VGGSound are the same as used in CAV-MAE (Gong et al., 2023). Additionally, to demonstrate the robustness of our model across different datasets, we perform further zero-shot retrieval experiments on MSR-VTT. Our model consistently outperforms existing methods pre-trained on AudioSet and achieves comparable results to models pre-trained on HowTo100M, such as Boggust (Boggust et al., 2019), Aranjelovic (Arandjelovic & Zisserman, 2017), and AVLnet (Noroozi & Favaro, 2016).
Table 3: Zero-shot audio-visual retrieval results on the MSR-VTT, AudioSet, and VGGSound. ${}^{\dagger}$esults reproduced on our environment. $*{}^{*}$odels pre-trained on the much larger dataset.
Video-to-Audio Audio-to-Video
Method R@1 R@5 R@10 R@1 R@5 R@10
MSR-VTT
$*{}^{*}$ 9.3 20.7 28.8 7.6 21.1 28.3
$*{}^{*}$ 11.9 25.9 34.7 12.6 26.3 33.7
$*{}^{*}$ 17.2 26.6 46.6 17.8 35.5 43.6
CAV-MAE 13.3 29.0 40.5 7.6 19.8 30.2
EquiAV (ours) 13.8 31.4 43.0 14.4 33.3 43.0
AudioSet
${}^{\dagger}$ 16.6 37.0 45.9 14.3 32.0 40.7
CAV-MAE 18.8 39.5 50.1 15.1 34.0 43.0
EquiAV (ours) 30.1 53.3 62.9 29.6 53.7 63.1
VGGSound
${}^{\dagger}$ 15.5 35.3 45.1 16.4 35.0 44.7
CAV-MAE 14.8 34.2 44.0 12.8 30.4 40.3
EquiAV (ours) 28.5 51.5 61.7 28.4 51.8 62.3
@@ 4.2 Ablation Studies
ZS Retrieval Fine-Tuning (mAP)
Intra. Inter. V2A A2V A V A-V
Inv. ($z_{a},z_{v}$) 25.2 24.3 40.1 22.9 42.3
Equi. ($z_{a},z_{v}$) 28.7 28.5 41.9 24.8 44.9
Equi. ($z^{\prime}_{a},z^{\prime}_{v}$) 28.1 27.5 41.5 24.2 43.9
Equi. ($\hat{z}_{a},\hat{z}_{v}$) 27.8 27.4 41.2 23.6 43.4
Equi. ($\bar{z}_{a},\bar{z}_{v}$) 30.1 29.6 42.4 25.7 46.6
@@
ZS Retrieval Fine-Tuning (mAP)
##\## representations V2A A2V A V A-V
1 27.8 27.4 41.2 23.6 43.4
4 28.3 29.0 41.4 24.4 44.9
8 29.4 29.2 41.9 24.9 45.5
16 30.1 29.6 42.4 25.7 46.6
@@
Similarity ZS Retrieval Fine-Tuning (mAP)
Method A V V2A A2V A V A-V
Linear 0.85 0.87 28.2 27.5 41.5 24.1 43.8
Hypernet. 0.87 0.90 29.6 29.6 42.0 25.2 44.5
MHA 0.90 0.92 30.1 29.6 42.4 25.7 46.6
ZS Retrieval Fine-Tuning (mAP)
Training Strategy V2A A2V A V A-V
Two-Stage 27.0 25.6 41.0 23.1 43.8
Alternating iterative 28.5 29.0 41.6 24.8 44.7
Joint 30.1 29.6 42.4 25.7 46.6
@@
(a) child singing
(b) playing electronic organ
(c) playing tympani
(d) children shouting
Figure 3: Qualitative results of EquiAV using VGGSound. Top Left: Original input image, Top Right: Zero-shot sound source localization result for the original input image. Bottom Left: Augmented image whose embedding is the closest to the centroid of equivariant embeddings, Botton Right: Zero-shot sound source localization result for the augmented image.
Zero-shot Retrieval Fine-Tuning (mAP)
Intra-modal Loss V2A A2V A V A-V
without pos (Eq. A) 21.9 21.0 39.5 22.0 42.6
with pos (Eq. 9) 30.1 29.6 42.4 25.7 46.6
@@
Table 8 shows the results of using different equivariant loss functions for intra-modal pairs. The first row excludes the positive pair in the denominator according to Equation A, while the second row includes it as in Equation 9. When Equation 9 is used for the intra-modal contrastive loss, the weight updates due to hard positives are relatively larger than those due to easy positives. In the context of equivariant contrastive learning, learning from hard positive with the transformation predictor is more effective, leading to better representation quality. Consequently, the semantically rich intra-modal representation promotes effective learning of audio-visual correspondence and audio-visual joint representations. The experimental results support our hypothesis. For further analytical insights, refer to Appendix A.

### v2

@@ Abstract
Recent advancements in self-supervised audio-visual representation learning have demonstrated its potential to capture rich and comprehensive representations. However, despite the advantages of data augmentation verified in many learning methods, audio-visual learning has struggled to fully harness these benefits, as augmentations can easily disrupt the correspondence between input pairs. To address this limitation, we introduce EquiAV, a novel framework that leverages equivariance for audio-visual contrastive learning. Our approach begins with extending equivariance to audio-visual learning, facilitated by a shared attention-based transformation predictor. It enables the aggregation of features from diverse augmentations into a representative embedding, providing robust supervision. Notably, this is achieved with minimal computational overhead. Extensive ablation studies and qualitative results verify the effectiveness of our method. EquiAV outperforms previous works across various audio-visual benchmarks. The code is available on https://github.com/JongSuk1/EquiAV
Machine Learning, ICML
@@ 1 Introduction
Audio and visual modalities play a pivotal role in how humans perceive their surroundings. Despite the differences in their characteristics, there exists an inherent correspondence between the two modalities. Learning such audio-visual correspondence from large-scale unlabeled video data in a self-supervised manner has recently become a major interest in the deep-learning research community. Among various approaches for audio-visual self-supervised learning, Audio-Visual Contrastive Learning has been popular due to its simplicity and effectiveness (Ma et al., 2021a, b; Morgado et al., 2021a, b; Patrick et al., 2021; Recasens et al., 2021; Wang et al., 2021; Sarkar & Etemad, 2023b).
@@
A key challenge in audio-visual contrastive learning is to enhance representational capability and diversity while maintaining the correspondence between two different modalities. The most straightforward and effective way to enrich the representation is to utilize data augmentation. However, data augmentation has not been extensively applied in the field of audio-visual contrastive learning. This is because correspondence between audio and visual modalities can be easily distorted due to data augmentations. For this reason, previous works utilize a very limited range of data augmentation and adopt alternatives to learn richer representations, such as incorporating masked data modeling (Gong et al., 2023; Huang et al., 2023) or utilizing temporal supervision (Sarkar & Etemad, 2023b; Jenni et al., 2023).
@@ 2 Related Works
Audio-visual contrastive learning has been one of the most popular approaches for learning the natural correspondence between audio and visual modalities, due to its simple intuition and powerful performance on the downstream tasks (Ma et al., 2021a, b; Morgado et al., 2021a, b; Patrick et al., 2021; Recasens et al., 2021; Wang et al., 2021; Owens & Efros, 2018; Sarkar & Etemad, 2023b). A common approach to audio-visual contrastive learning is to learn the context of the synchronous relationship between audio and visual inputs (Korbar et al., 2016; Alwassel et al., 2020a; Morgado et al., 2021b; Sarkar & Etemad, 2023b). On the other hand, several works (Georgescu et al., 2023; Haliassos et al., 2022) adopt masked modeling techniques that are designed to reconstruct the masked raw inputs or predict the masked context features. Recently, CAV-MAE (Gong et al., 2023) and MAViL (Huang et al., 2023) have combined contrastive learning and masked data modeling techniques to learn complementary representations.
@@ 3.2 EquiAV
In our framework, typical visual-based augmentations are employed for visual inputs. However, unlike the abundant augmentations available for an image or video, the options for the audio modality are relatively limited. To address this issue, we convert the audio signal into a spectrogram, followed by the application of visual-based augmentations in the same manner. The augmentation information is encoded into real vectors, denoted as $\mathbf{t}_{a}$ and $\mathbf{t}_{v}$. These augmentation vectors parameterize how much each augmentation is applied to the input data. A detailed explanation of data augmentation parameterization is provided in Appendix D.
@@
$h^{\text{rep}}=\mathbb{E}_{h^{\prime}\sim D_{h}}\left[h^{\prime}\right].$ (11)
The vector $h^{\text{rep}}$ integrates information from multiple augmented data, securing more general features while also providing robustness against undesirable effects of augmentation. However, it is not feasible to generate all transformations and obtain the representative vector $h^{\text{rep}}$, we aim to approximate it with the centroid of sampled equivariant representations.
$\displaystyle h^{\text{rep}}_{m}\approx\bar{h}_{m}$ $=\frac{1}{S}\sum_{i=1}^{S}\hat{h}_{m}^{i}=\frac{1}{S}\sum_{i=1}^{S}u_{m}(h_{m}^{i},\mathbf{t}_{m}^{i}).$ (12)
@@
    $\mathcal{L}^{\text{inter}}=\frac{1}{N}\sum_{i=1}^{N}\ell^{\text{inter}}(i)$
@@ 4 Experiments
Table 1: Audio-visual event classification performance on AudioSet and VGGSound. A: Audio-only, V: Visual-only, A-V: Audio-visual. IN SL: ImageNet supervised learning, SSL: Self-supervised learning, Non-standard train/test split.
Pretrain AudioSet-20K (mAP) AudioSet-2M (mAP) VGGSound (acc.) Method A V A-V A V A-V A V A-V GBlend (Wang et al., 2020) - 29.1 22.1 37.8 32.4 18.8 41.8 - - - Perceiver (Jaegle et al., 2021) - - - - 38.4 25.8 44.2 - - - Attn AV (Fayek & Kumar, 2021) IN SL - - - 38.4 25.7 46.2 - - - MBT (Nagrani et al., 2021) IN21K SL 31.3 27.7 43.9 41.5 31.3 49.6 52.3 51.2 64.1 CAV-MAE (Gong et al., 2023) SSL 37.7 19.8 42.0 46.6 26.2 51.2 59.5 47.0 65.5 AudiovisualMAE (Georgescu et al., 2023) SSL - - - 46.6 31.1 51.8 57.2 50.3 65.0 MAViL (Huang et al., 2023) SSL 41.8 24.8 44.9 48.7 30.3 53.3 60.8 50.9 67.1 EquiAV (ours) SSL 42.4 25.7 46.6 49.1 30.1 54.6 61.0 50.7 67.1
The model is pre-trained on the AudioSet-2M in a self-supervised manner, without the use of labels. The audio encoder $f_{a}$ and visual encoder $f_{v}$ are initialized with the self-supervised pre-trained ViT-B/16 model of MAE (He et al., 2022). We employ 3-layer MLP with layer normalization for intra-modal and inter-modal projection heads. Note that the multi-head attention layer and the feed-forward layer of the transformation predictor are similar to those used in Transformers (Vaswani et al., 2017; Chen et al., 2021), which include layer normalization layer and residual connection. We set $\lambda^{\text{inter}}$, $\lambda^{\text{inter}}_{a}$ and $\lambda^{\text{inter}}_{v}$ all the same as 1. More detailed experimental settings are explained in Appendix D and Appendix E.
Table 2: Audio classification accuracy of finetuning on ESC-50, SPC-v1, and FSD-50K. Results reproduced on our environment. [^2] Weights from https://github.com/YuanGongND/cav-mae
Method ESC-50 SPC-v1 FSD-50K CAV-MAE-Scale++ (Gong et al., 2023) 83.2 97.1 45.5 MAViL (Huang et al., 2023) 94.4 97.4 - XKD (Sarkar & Etemad, 2024) 96.5 - 58.5 EquiAV (ours) 96.0 97.8 62.6
Table 3: Action recognition accuracy of transfer learning on UCF101, HMDB51, and Kinetics400. High temporal resolution with industry-level computational resources.
Method UCF101 HMDB51 Kinetics400 XDC (Alwassel et al., 2020b) 85.3 56.0 - MMV (Alayrac et al., 2020) 83.9 60.0 - BraVe (Recasens et al., 2021) 90.0 63.6 - AVID (Morgado et al., 2021c) - - 48.9 CrissCross (Sarkar & Etemad, 2023a) 87.7 56.2 50.1 XKD (Sarkar & Etemad, 2024) 88.4 62.2 56.5 EquiAV (ours) 89.7 64.4 57.3
@@ 4.1 Main Results
Comparison on the Single-modal Benchmarks.
To demonstrate the performance of our framework on the single-modal downstream tasks, we fine-tune the audio branch of EquiAV on audio classification datasets, including the ESC-50 (Piczak, 2015), SPC-v1 (Warden, 2018), and FSD-50K (Fonseca et al., 2021). According to the results reported in Table 2, our method shows superior performance compared to other self-supervised learning-based methods. We also evaluate the video branch of EquiAV on various action recognition datasets, including UCF101 (Soomro et al., 2012), HMDB51 (Kuehne et al., 2011), and Kinetics400 (Kay et al., 2017). As demonstrated in Table 3, our method surpasses previous SOTA audio-visual SSL methods on uni-modal visual benchmark datasets. The comparison is based on results from other audio-visual SSL methods that also utilize the same pre-training dataset, AudioSet-2M.
Table 4: Zero-shot audio-visual retrieval results on the MSR-VTT, AudioSet, and VGGSound. Results reproduced on our environment. Models pre-trained on the much larger dataset.
Video-to-Audio Audio-to-Video Method R@1 R@5 R@10 R@1 R@5 R@10 MSR-VTT Boggust 9.3 20.7 28.8 7.6 21.1 28.3 Aranjelovic 11.9 25.9 34.7 12.6 26.3 33.7 AVLnet 17.2 26.6 46.6 17.8 35.5 43.6 CAV-MAE 13.3 29.0 40.5 7.6 19.8 30.2 EquiAV (ours) 13.8 31.4 43.0 14.4 33.3 43.0 AudioSet CAV-MAE-Scale++ 16.6 37.0 45.9 14.3 32.0 40.7 CAV-MAE 18.8 39.5 50.1 15.1 34.0 43.0 EquiAV (ours) 30.1 53.3 62.9 29.6 53.7 63.1 VGGSound CAV-MAE-Scale++ 15.5 35.3 45.1 16.4 35.0 44.7 CAV-MAE 14.8 34.2 44.0 12.8 30.4 40.3 EquiAV (ours) 28.5 51.5 61.7 28.4 51.8 62.3
@@
Table 4 reports the performance of zero-shot retrieval on audio-visual datasets, based on the similarity between the audio-visual embeddings. The sample lists used in the experiments for AudioSet and VGGSound are the same as those used in CAV-MAE (Gong et al., 2023). Additionally, to demonstrate the robustness of our model across different datasets, we perform further zero-shot retrieval experiments on MSR-VTT. Our model consistently outperforms existing methods pre-trained on AudioSet and achieves comparable results to models pre-trained on HowTo100M, such as Boggust (Boggust et al., 2019), Aranjelovic (Arandjelovic & Zisserman, 2017), and AVLnet (Noroozi & Favaro, 2016).
ZS Retrieval Fine-Tuning (mAP) Intra. Inter. V2A A2V A V A-V Inv. ($z_{a},z_{v}$) 25.2 24.3 40.1 22.9 42.3 Equi. ($z_{a},z_{v}$) 28.7 28.5 41.9 24.8 44.9 Equi. ($z^{\prime}_{a},z^{\prime}_{v}$) 28.1 27.5 41.5 24.2 43.9 Equi. ($\hat{z}_{a},\hat{z}_{v}$) 27.8 27.4 41.2 23.6 43.4 Equi. ($\bar{z}_{a},\bar{z}_{v}$) 30.1 29.6 42.4 25.7 46.6
@@ 4.2 Ablation Studies
ZS Retrieval Fine-Tuning (mAP) ##\## representations V2A A2V A V A-V 1 27.8 27.4 41.2 23.6 43.4 4 28.3 29.0 41.4 24.4 44.9 8 29.4 29.2 41.9 24.9 45.5 16 30.1 29.6 42.4 25.7 46.6
Similarity ZS Retrieval Fine-Tuning (mAP) Method A V V2A A2V A V A-V Linear 0.85 0.87 28.2 27.5 41.5 24.1 43.8 Hypernet. 0.87 0.90 29.6 29.6 42.0 25.2 44.5 MHA 0.90 0.92 30.1 29.6 42.4 25.7 46.6
@@
ZS Retrieval Fine-Tuning (mAP) Training Strategy V2A A2V A V A-V Two-Stage 27.0 25.6 41.0 23.1 43.8 Alternating iterative 28.5 29.0 41.6 24.8 44.7 Joint 30.1 29.6 42.4 25.7 46.6
@@
Zero-shot Retrieval Fine-Tuning (mAP) Intra-modal Loss V2A A2V A V A-V without pos (Eq. A) 21.9 21.0 39.5 22.0 42.6 with pos (Eq. 9) 30.1 29.6 42.4 25.7 46.6
@@
Table 9 shows the results of using different equivariant loss functions for intra-modal pairs. The first row excludes the positive pair in the denominator according to Equation A, while the second row includes it as in Equation 9. When Equation 9 is used for the intra-modal contrastive loss, the weight updates due to hard positives are relatively larger than those due to easy positives. In the context of equivariant contrastive learning, learning from hard positive with the transformation predictor is more effective, leading to better representation quality. Consequently, the semantically rich intra-modal representation promotes effective learning of audio-visual correspondence and audio-visual joint representations. The experimental results support our hypothesis. For further analytical insights, refer to Appendix C.